OpenStudy (anonymous):

Find three numbers whose sum is 6 and whose sum of squares is a minimum. Can someone help me set this up?

OpenStudy (bahrom7893):

x+y+z=6 x^2+y^2+z^2 is a minimum.

OpenStudy (bahrom7893):

hmm u sure this is algebra 2?

OpenStudy (anonymous):

I said above algebra ;)

OpenStudy (bahrom7893):

oh u just gotta set this up.

OpenStudy (anonymous):

I have to solve using partials, but I just can't get it set up right

OpenStudy (bahrom7893):

u learn partial derivatives in algebra?

OpenStudy (anonymous):

I'm in calc III

OpenStudy (bahrom7893):

lol sorry man, look up lagrange multipliers, I totally forgot my calc 3. @Zarkon can u help this guy out?

OpenStudy (anonymous):

thanks anyway :)

OpenStudy (zarkon):

yes...this problem is trivial with lagrange multipliers

OpenStudy (zarkon):

just compute \[\nabla(x^2+y^2+z^2)=\lambda \nabla(x+y+z-6)\]

OpenStudy (anonymous):

i got it. it is 2,2,2 i was not expanding my z. thanks guys!

OpenStudy (zarkon):

that's it. though I'm not sure what you mean about expanding your z \[2x=\lambda\] \[2y=\lambda\] \[2z=\lambda\] \[\Rightarrow x=y=z\] \[x+y+z=6\Rightarrow x+x+x=6\Rightarrow x=2\Rightarrow y=2\Rightarrow z=2\]